# Designing Library of Skill-Agents for Hardware-Level Reusability

### 04 Mar 2024 v1

@@ A.1 OP-PR and PR-OP
A typical example of an object in the OP (M=0, D=1, C=2) state is a peg fully inserted into a hole with touching the bottom of the hole. The possible motion of the peg is limited to the motion of pulling it out from the hole. When the peg is slightly pulled out, it detaches from the bottom surface. In other words, the transition occurs from the detachment d-state to the maintenance d-state. A3 is applicable to this direction. Two dimensions orthogonal to the motion direction remain constrained both before and after the transition, requiring control to maintain the constraint d-state throughout the transition, as indicated by B3.
@@
  Reward OP-PR (PTG31 (Drawer-open) skill)
@@
  Reward PR-OP (PTG33 (Drawer-close) skill)
@@ A.2 PR-NC and NC-PR
    if S = goal-s AND T = goal-t AND
       U = goal-u AND F-u < delta-zero AND
       F-t < delta-zero, then reward
@@
The transition from NC (M=3, D=0, C=0) to PR (M=1, D=0, C=2) corresponds to the opposite scenario of the previous example, where a peg in mid-air is suddenly inserted into a hole. Concerning the direction of motion, the maintenance d-state is maintained. In other words, A1 is applicable, and the terminal condition is defined solely based on positional information, given by the demonstration. Regarding the directions orthogonal to the motion, both directions are unconstrained before the transition, and after the transition, they transition into the constrained d-state due to environmental constraints. Specifically, B9 is applicable, indicating that visual feedback in those dimensions is necessary.
@@ A.3 PR-PC and PC-PR
An example of a transition from PR (M=1, D=0, C=2) to PC (M=2, D=1, C=0) involves pulling a peg out from within a hole. As the peg moves, one direction orthogonal to the motion direction maintains contact with the continuous surface, such as the bottom of the hole, even after leaving the hole, while the other direction becomes unconstrained and enters the maintenance d-state. In this case, the detachment d-state occurs in one direction (with the bottom surface), and the maintenance d-state is reached in the other.
Concerning the direction of motion, it remains in the maintenance d-state, making A1 applicable. On the other hand, in dimensions orthogonal to the motion, for one dimension, there is a transition from the constraint d-state to the maintenance d-state (B8 is applicable), and for another dimension, there is a transition from the constraint d-state to the detachment d-state (B7 is applicable).
@@
    if S = goal-s AND T = goal-t AND
       F-t < delta-zero, then reward
@@
The direction of motion, both before and after the transition, maintains the maintenance d-state, making A1 applicable. In one dimension orthogonal to the motion, where the detachment d-state transits to the constraint d-state, B6 is applicable. Namely, by maintaining the detachment d-state, the system automatically enters the constraint d-state. For the other dimension, transiting from the maintenance d-state to the constraint d-state, B9 is applicable, indicating the need for visual feedback.
@@ A.4 PR-TR and TR-PR
    if S = goal-s AND T = goal-t AND
       F-t < delta-zero, then reward
@@ A.5 PR-OT and OT-PR
The transition form PR (M=1, D=0, C=2) to OT (M=1, D=1, C=1) involves the transition from the constraint d-state to the detachment d-state in one of the two constraint d-state dimensions. For example, when a peg in a hole is pulled out, a portion of one set of opposing constraint surfaces is removed, resulting in a detachment d-state, while the other dimension remains in a constraint d-state.
Regarding the direction of motion, it remains in the maintenance d-state, and A1 can be applied. As for the second dimension to maintain the constraint d-state, B3 can be applied: The third dimension transiting from the constraint d-state to the detachment d-state, and B7 can be applied.
@@
The transition from OT (M=1, D=1, C=1) to PR (M=1, D=0, C=2) involves the dimension that was in the detachment d-state transiting to the constraint d-state. For this dimension, maintaining contact alone is sufficient to naturally transition from the detachment d-state to the constraint d-state.
As for the direction of motion, the maintenance d-state remains throughout the transition, and A1 can be applied. One dimension orthogonal to the motion remains in the constraint d-state, and B3 can be applied. In the other orthogonal dimension to the motion, a transition occurs from the detachment d-state to the constraint d-state. B6 is applicable.
@@ A.6 OT-NC and NC-OT
In such case, concerning the direction of motion, the maintenance d-state is maintained. Therefore, A1 can be applied. In one dimension orthogonal to the motion, a transition occurs from the detachment d-state to the maintenance d-state. Therefore, B5 can be applied. In the other orthogonal dimension to the motion, the constraint d-state transits to the maintenance d-state. Therefore, B8 is applicable.
@@
      if S = goal-s AND T = goal-t AND
         F-t < delta-zero AND U = goal-u AND
         F-u < delta-zero, then reward
@@
As for the motion direction, the maintenance d-state is maintained, and A1 can be applied. In one dimension orthogonal to the motion, there is a transition from the maintenance d-state to the detachment d-state. A visual sensor is required for this transition. In the other dimension orthogonal to the motion, there is a transition from the maintenance d-state to the constraint d-state, and this also requires a visual sensor.
@@ A.7 OT-PC and PC-OT
    if S = goal-s AND U = goal-u AND
       F-u < delta-zero, then reward
@@
      if S = goal-s AND T = goal-t
         F-t < delta-zero, then reward
@@
As for the transitions from PC (M=2, D=1, C=0) to OT (M=1, D=1, C=1), there are also two scenarios. In the PC-OT-a case, regarding to the direction of motion, the maintenance d-state is maintained. A1 can be applied. In one orthogonal direction to the motion, the detachment d-state is maintained. B2 can be applied. In the other orthogonal direction, the maintenance d-state transits to the constraint d-state and B9 can be applied.
@@ A.8 OT-TR and TR-OT
    if S = goal-s AND T = goal-t AND
       F+s < delta-zero, then reward
@@
      if S = goal-s AND U = goal-u AND
         F-u < delta-zero, then reward
@@
    if T = goal-t AND F-s > delta-zero,
@@ B.1 TR-TR
Regarding the transition from TR (M=2, D=0, C=1) to TR, in the direction of motion, the maintenance d-state is maintained. A1 can be applied. In one orthogonal direction to the motion, the maintenance d-state is maintained. B1 can be applied. In the other orthogonal direction, the constraint d-state is maintained. B2 can be applied.
@@ B.2 OT-OT
The OT state consists of a set of two Kuhn-Tucker solution classes. OT1 (M=1, D=1, C=1)) has the solution domain on a semi-circular arc on the great circle, while OT2 (M=0, D=2, C=1) has a partial arc of the great circle as the domain of solutions.
@@
In the OT1 state, there are two possible directions of motion: motion in the detachment direction and motion along the detachment surface. However, to transit to OT1, motion along the detachment surface is required. In this case, the d-state of motion direction is such that the maintenance d-state is preserved. A1 can be applied. On the other hand, in one orthogonal direction to the motion, the detachment d-state is maintained. B2 can be applied. In the other orthogonal dimension, the constraint d-state is maintained. B3 can be applied.
@@
The region of solutions for OT1, with further constraints, becomes a partial arc, resulting in OT2. Therefore, the transition form OT1 (M=2, D=0, C=1) to OT2 (M=1, D=1, C=1) occurs in cases where motion along the detachment surface results in encountering another contact surface. Consequently, concerning the direction of motion, the maintenance d-state transits to the detachment d-state. A2 can be applied. One orthogonal direction to the motion maintains the detachment d-state. B2 can be applied. The other orthogonal direction to the motion maintains the constraint d-state. B3 can be applied.
@@
The transition from OT2 (M=1, D=1, C=1) to OT1 (M=2, D=0, C=1) is the reverse of the previous transition, where the solution domain restricted to a partial arc transits, through detachment motion, to a semi-circle region. Therefore, concerning the direction of motion, a transition from the detachment d-state to a maintenance d-state occurs. Therefore A3 can be applied. One orthogonal direction to the motion maintains the detachment d-state. B2 can be applied. The other orthogonal direction to the motion maintains the constraint d-state. B3 can be applied.

### 20 Mar 2024 v2

@@ A.1 OP-PR and PR-OP
A typical example of an object in OP (M=0, D=1, C=2) is a peg fully inserted into a hole with touching the bottom of the hole. The possible motion of the peg is limited to the motion of pulling it out from the hole. When the peg is slightly pulled out, it detaches from the bottom surface. In other words, the transition occurs from the detachment d-state to the maintenance d-state. A3 is applicable to this direction. Two dimensions orthogonal to the motion direction remain constrained both before and after the transition, requiring control to maintain the constraint d-state throughout the transition, as indicated by B3.
@@
  Reward OP-PR (PTG31 (Drawer-open) task)
@@
  Reward PR-OP (PTG33 (Drawer-close) task)
@@ A.2 PR-NC and NC-PR
    if S = goal-s AND F-t < delta-zero AND
       T = goal-t AND F-u < delta-zero AND
       U = goal-u, then reward
@@
The transition from NC (M=3, D=0, C=0) to PR (M=1, D=0, C=2) corresponds to the opposite scenario of the previous example, where a peg in the air is suddenly inserted into a hole. Concerning the direction of motion, the maintenance d-state is maintained. In other words, A1 is applicable, and the terminal condition is defined solely based on positional information, given by the demonstration. Regarding the directions orthogonal to the motion, both directions are unconstrained before the transition, and after the transition, they transit into the constrained d-state due to environmental constraints. Specifically, B9 is applicable, indicating that visual feedback in those dimensions is necessary.
@@ A.3 PR-PC and PC-PR
An example of a transition from PR (M=1, D=0, C=2) to PC (M=2, D=1, C=0) involves pulling a peg out from within a hole. As the peg moves, one direction orthogonal to the motion direction maintains contact with the continuous surface even after leaving the hole, while the other direction becomes unconstrained and enters the maintenance d-state. In this case, the detachment d-state occurs in one direction (with the bottom surface), and the maintenance d-state is reached in the other.
Concerning the direction of motion, it remains in the maintenance d-state, making A1 applicable. On the other hand, in dimensions orthogonal to the motion, for one dimension, there is a transition from the constraint d-state to the maintenance d-state (B8 is applicable), and for the other dimension, there is a transition from the constraint d-state to the detachment d-state (B7 is applicable).
@@
    if S = goal-s AND F-t < delta-zero AND
       T = goal-t, then reward
@@
The direction of motion, both before and after the transition, maintains the maintenance d-state, making A1 applicable. In one dimension orthogonal to the motion, where the detachment d-state transits to the constraint d-state, B6 is applicable. Namely, by maintaining the detachment d-state, the system automatically enters the constraint d-state. For the other dimension, the transition from the maintenance d-state to the constraint d-state occurs and B9 is applicable, indicating the need for visual feedback.
@@ A.4 PR-TR and TR-PR
    if S = goal-s AND F-t < delta-zero AND
       T = goal-t, then reward
@@ A.5 PR-OT and OT-PR
The transition from PR (M=1, D=0, C=2) to OT (M=1, D=1, C=1) involves the transition from the constraint d-state to the detachment d-state in one of the two constraint d-state dimensions. For example, when a peg in a hole is pulled out, a portion of one set of opposing constraint surfaces is removed, resulting in a detachment d-state, while the other dimension remains in a constraint d-state.
Regarding the direction of motion, it remains in the maintenance d-state, and A1 can be applied. As for the second dimension to maintain the constraint d-state, B3 can be applied. The third dimension transits from the constraint d-state to the detachment d-state and B7 can be applied.
@@
The transition from OT (M=1, D=1, C=1) to PR (M=1, D=0, C=2) involves the dimension that was in the detachment d-state transiting to the constraint d-state. For this dimension, maintaining contact alone is sufficient to naturally transit from the detachment d-state to the constraint d-state.
As for the direction of motion, the maintenance d-state remains throughout the transition and A1 can be applied. One dimension orthogonal to the motion remains in the constraint d-state and B3 can be applied. In the other orthogonal dimension to the motion, a transition occurs from the detachment d-state to the constraint d-state. B6 is applicable.
@@ A.6 OT-NC and NC-OT
In such case, concerning the direction of motion, the maintenance d-state is maintained. Therefore A1 can be applied. In one dimension orthogonal to the motion, a transition occurs from the detachment d-state to the maintenance d-state. Therefore, B5 can be applied. In the other orthogonal dimension to the motion, the constraint d-state transits to the maintenance d-state. Therefore, B8 is applicable.
@@
      if S = goal-s AND F-t < delta-zero AND
         T = goal-t AND F-u < delta-zero AND
         U = goal-u, then reward
@@
As for the motion direction, the maintenance d-state is maintained, and A1 can be applied. In one dimension orthogonal to the motion, there is a transition from the maintenance d-state to the detachment d-state and B4 can be applied; a visual sensor is required for this transition. In the other dimension orthogonal to the motion, there is a transition from the maintenance d-state to the constraint d-state and B9 can be applied. This also requires a visual sensor.
@@ A.7 OT-PC and PC-OT
    if S = goal-s AND F-u < delta-zero AND
       U = goal-u, then reward
@@
      if S = goal-s AND F-t < delta-zero AND
         T = goal-t, then reward
@@
As for the transition from PC (M=2, D=1, C=0) to OT (M=1, D=1, C=1), there are also two scenarios. In the PC-OT-a case, regarding to the direction of motion, the maintenance d-state is maintained. A1 can be applied. In one orthogonal direction to the motion, the detachment d-state is maintained. B2 can be applied. In the other orthogonal direction, the maintenance d-state transits to the constraint d-state and B9 can be applied.
@@ A.8 OT-TR and TR-OT
    if F+s < delta-zero AND S = goal-s AND
       T = goal-t, then reward
@@
      if S = goal-s AND F-u < delta-zero AND
         U = goal-u, then reward
@@
    if F-s > delta-zero AND T = goal-t,
@@ B.1 TR-TR
Regarding the transition from TR (M=2, D=0, C=1) to TR, in the direction of motion, the maintenance d-state is maintained. A1 can be applied. In one orthogonal direction to the motion, the maintenance d-state is maintained. B1 can be applied. In the other orthogonal direction, the constraint d-state is maintained. B3 can be applied.
@@ B.2 OT-OT
The OT state consists of a set of two Kuhn-Tucker solution classes. OT1 (M=1, D=1, C=1) has the solution domain on a semi-circular arc on the great circle, while OT2 (M=0, D=2, C=1) has a partial arc of the great circle as the domain of solutions.
@@
In OT1, there are two possible directions of motion: motion in the detachment direction and motion along the detachment surface. However, to transit to OT1, motion along the detachment surface is required. In this case, the d-state of motion direction is such that the maintenance d-state is preserved. A1 can be applied. On the other hand, in one orthogonal direction to the motion, the detachment d-state is maintained. B2 can be applied. In the other orthogonal dimension, the constraint d-state is maintained. B3 can be applied.
@@
The region of solutions for OT1, with further constraints, becomes a partial arc, resulting in OT2. Therefore, the transition from OT1 (M=1, D=1, C=1) to OT2 (M=0, D=2, C=1) occurs in cases where motion along the detachment surface results in encountering another contact surface. Consequently, concerning the direction of motion, the maintenance d-state transits to the detachment d-state. A2 can be applied. One orthogonal direction to the motion maintains the detachment d-state. B2 can be applied. The other orthogonal direction to the motion maintains the constraint d-state. B3 can be applied.
@@
The transition from OT2 (M=0, D=2, C=1) to OT1 (M=1, D=1, C=1) is the reverse of the previous transition, where the solution domain restricted to a partial arc, transits through detachment motion to a semi-circle region. Therefore, concerning the direction of motion, a transition from the detachment d-state to a maintenance d-state occurs. Therefore A3 can be applied. One orthogonal direction to the motion maintains the detachment d-state. B2 can be applied. The other orthogonal direction to the motion maintains the constraint d-state. B3 can be applied.